# Unexpected Effects of Disorder on Current Fluctuations in the Symmetric Simple Exclusion Process

### v1

@@ I Introduction
Single file diffusion, where particles diffuse in a narrow channel and cannot pass through each other, is a pedagogical model for interacting many-body systems. It is well known that, on the infinite line, the mean square displacement (MSD) of the tracer particle is given by $\Braket{\delta x(t)^{2}}\propto t^{1/2}$ in the long time limit Harris (1965), where $\delta x(t)$ is the displacement of the tracer particle and $\Braket{\cdot}$ is the ensemble average. Various transport phenomena in narrow channels show single file diffusion, for example, molecular diffusion in zeolites Hahn et al. (1996); Kukla et al. (1996), DNA-binding proteins Li et al. (2009), transport of confined colloidal particles Wei et al. (2000); Lutz et al. (2004), and transport of water molecules in carbon nanotubes Berezhkovskii and Hummer (2002).
The symmetric simple exclusion process (SSEP) is a fundamental model of single-file diffusion Spitzer (1970), where particles with hard-core interactions move on a one-dimensional lattice. With its simplicity, this model has led to the derivation of mathematically rigorous results. On the infinite lattice, the large deviation function of the integrated current through the origin has been computed independently by Bethe ansatz Derrida and Gerschenfeld (2009) and the macroscopic fluctuation theory Mallick et al. (2022). Additionally, the large deviation function of the tracer position has been derived Imamura et al. (2017, 2021). The generalized density profile, which is the correlation between the observation (for example, the tracer position, the integrated current through the origin, and the generalized current) and the density, has also been derived Poncet et al. (2021); Grabsch et al. (2022, 2023). Under the periodic boundary condition, the large deviation function of the integrated current has also been obtained Appert-Rolland et al. (2008). In particular, the second moment of the total integrated current $Q(t)$ in time $t$ is derived as a function of the particle density $\rho$:
@@
where $\tau$ is the mean waiting time, and $L$ is the system size Appert-Rolland et al. (2008).
@@
Single-particle diffusion in a heterogeneous environment shows various anomalous behaviors. Such environments are represented by random energy landscapes. The random energy landscape can be categorized into two types: an annealed energy landscape, which changes with time, and a quenched energy landscape, which remains static. The random walk on an annealed energy landscape is often modeled by the continuous-time random walk. In this model, waiting times are independent and identically distributed (IID) random variables, and the MSD exhibits anomalous diffusion when the mean waiting time diverges Metzler and Klafter (2000). The random walk on a quenched energy landscape is called the quenched trap model Bouchaud and Georges (1990). In the infinite system, the MSD shows anomalous diffusion when the mean waiting time diverges Bouchaud and Georges (1990). In the finite system, observables such as the drift, the diffusion coefficient, and the first passage time exhibit sample-to-sample fluctuations, resulting in non-self-averaging behaviors Akimoto et al. (2016); *AkimotoBarkaiSaito2018; Luo and Yi (2018); Akimoto and Saito (2019, 2020).
@@
The influence of heterogeneity on the single-file diffusion has garnered significant attention in the field of statistical mechanics Sorkin and Dean (2023); Metzler et al. (2014); Sanders et al. (2014); Chou (1998); Ben-Naim and Krapivsky (2009). Recently, the single file diffusion in the system with periodically varying potential and diffusivity has been studied, and the MSD of the tracer particle has been obtained Sorkin and Dean (2023). Specifically, the formula of the MSD is the same as that for the normal single-file diffusion. However, the diffusion coefficient is replaced with the effective diffusion coefficient for a single particle diffusing in the system with periodically varying potential and diffusivity. For the SSEP on an annealed energy landscape, when the mean or variance of waiting time diverges, the MSD is $\Braket{\delta x(t)^{2}}\propto\log^{1/2}t$ or $\Braket{\delta x(t)^{2}}\propto t^{\gamma}$ with $20<\gamma<1/$ 2, respectively Metzler et al. (2014); Sanders et al. (2014), i.e., the diffusion in the SSEP on an annealed energy landscape is slower than in the normal SSEP. The impacts of disorder in the asymmetric simple exclusion process, where particles in the SSEP are influenced by an external field like an electric field, have been a subject of research for many years Tripathy and Barma (1998); Enaud and Derrida (2004); Harris and Stinchcombe (2004); Juhász et al. (2006); Nossan (2013); Concannon and Blythe (2014); Bahadoran and Bodineau (2015); Banerjee and Basu (2020); Sakai and Akimoto (2023a); *PhysRevE.107.054131. While several effects resulting from the interplay between a heterogeneous environment and a many-body system have been uncovered, the study of the SSEP on a quenched random energy landscape remains unexplored. Such heterogeneity is prevalent in various diffusion phenomena, such as the movement of proteins on DNA Granéli et al. (2006); Wang et al. (2006) and water transportation in aquaporin Yamamoto et al. (2014) and OmpF van Hijkoop et al. (2007). Therefore, understanding current fluctuations in the SSEP on a quenched random energy landscape is crucial.
@@ II Model
We study the SSEP with a defect site to unravel the effects of heterogeneity of the environment. The system consists of $N$ particles on a ring of $L$ sites, where each site can accommodate at most one particle. The defect site has a longer mean waiting time compared to other sites, inhibiting particle movement. The mean waiting times at site $i$ are defined as follows:
@@
We explicitly describe the dynamics of particles. Initially, the particles are in an equilibrium ensemble of configurations. The equilibrium ensemble is given by the configuration after particles are randomly placed and diffuse for a long time. In infinite systems, effects of the initail ensemble are everlasting Leibovich and Barkai (2013). At time $t=0$, each particle $i$ is assigned the waiting time $t_{i}$ from $\psi^{(x_{i})}(t_{i})$, where $x_{i}$ is the site number of particle $i$. Once the waiting time $t_{i}$ has elapsed, particle $i$ attempts to jump to the right site with a probability of $1/2$ and to the left site with a probability of $1/2$. The jump is successful only if the target site is unoccupied. Following a successful or unsuccessful jump attempt, particle $i$ is given a new waiting time based on either $\psi^{(x_{i}\pm 1)}(t_{i})$ or $\psi^{(x_{i})}(t_{i})$, respectively.
@@ III.1 Correlation between total integrated current and occupation number
To obtain the correlation between the total integrated current and the occupation number, we derive $\Braket{Q\eta_{2}}_{L-1}$. When $N=L-1$, there is a single hole. The hole at site $i$ jump to the right site with rate $1/(2\tau_{i+1})$ and to the left site with rate $1/(2\tau_{i-1})$. Using the result in Ref. Derrida (1983), the second moment of the total integrated current for $N=L-1$ is given by
@@ III.2 Second moment of total integrated current
Based on the results of the previous subsection, the second moment of the total integrated current is represented by
@@
We note that the function $g(N,L,\mu)$ follows Eq. (30), and $g(2,L,\mu)$ and $g(L-2,L,\mu)$ are obtained by the numerical simulations. We compare the theoretical result, Eq. (33), with the numerical result. Figure 1 shows good agreement between numerical simulations and the theory. The density dependence of $\kappa(N)$ for the SSEP with a defect site is different from that for the homogeneous system (1). In particular, it becomes asymmetric around $\rho=1/2$. Furthermore, in the intermediate and high-density regime, $\kappa(N)$ for the SSEP with a defect site is smaller than that for the homogeneous system. This result is that the defect hiders the particle movement, leading to a reduction in fluctuations. On the other hand, in the low-density regime, $\kappa(N)$ for the SSEP with a defect site exceeds that for the homogeneous system. This property is counterintuitive because despite the defect impeding the movement of particles, the fluctuations increase. In terms of diffusivity, adjusting the defect allows for control over the particle’s diffusivity.
@@ IV Application: SSEP on a quenched random energy landscape
Figure 2: Relation between $\kappa(N)$ and density, corresponding to the SSEP on a quenched random energy landscape ($L=100$, $T=1$, $\tau_{c}=1$): (a)-(d) $\Braket{E}<T$ and (e)-(h) $\Braket{E}>T$. (a),(e) The energy depth distribution is the exponential distribution ($T/T_{g}=1.5$ for (a) and $T/T_{g}=0.5$ for (e)). (b),(f) The energy depth distribution is the gamma distribution ($T/T_{g}=2.5,~{}k=2$ for (b) and $T/T_{g}=2.5,~{}k=4$ for (f)). (c),(g) The energy depth distribution is the half-gaussian distribution ($\sigma^{2}=1$ for (c) and $\sigma^{2}=5$ for (g)). (d),(h) The energy depth distribution is the uniform distribution ($\varepsilon=1$ for (d) and $\varepsilon=3$ for (h)). Circles are the numerical simulation of the dynamics of the SSEP on a quenched random energy landscape. Solid lines represent Eq. (43), where we use $g(2,L,\mu)$ and $g(L-2,L,\mu)$ obtained by the numerical simulations of the dynamics of the SSEP with a defect site.
Here, we apply our theory of the current fluctuations of the SSEP with a defect site to the SSEP on a quenched random energy landscape. Quenched disorder means that when realizing the random energy landscape, it does not change with time. At each site, the depth $E>0$ of the energy trap is randomly assigned. The energies of the traps are IID random variables following distribution $\phi(E)$. A particle escapes from the trap and attempts to jump to the nearest neighbors. Waiting times at site $i$ are IID random variables with an exponential distribution, $\psi^{(i)}(\tau)=\tau_{i}^{-1}\exp(-\tau/\tau_{i})$. The mean waiting time follows the Arrhenius law, $\tau_{i}=\tau_{c}\exp(E_{i}/T)$, where $E_{i}$ is the depth of the energy at site $i$, $T$ the temperature, and $\tau_{c}$ typical time.
@@
Based on the Arrhenius law, the probability which the mean waiting time $\tau$ is smaller than $x$ is given by $\Pr(\tau\leq x)\cong\Pr(E\leq T\ln(x/\tau_{c}))$. Hence, the distributions of the mean waiting times follow
@@
$\psi(\tau)\cong\frac{\alpha^{n}\tau_{c}^{\alpha}}{\Gamma(n)}\left[\ln\left(\frac{\tau}{\tau_{c}}\right)\right]^{n-1}\tau^{-(1+\alpha)}\quad(\tau\geq\tau_{c})$ (39)
@@ V Discussion
We compare the current fluctuation in the single-particle system and the many-particle system. In the single-particle system, the current fluctuation is the same as the mean square displacement, i.e., $\kappa(1)=\bar{\tau}^{-1}$, where $\bar{\tau}=\sum_{i=1}^{L}\tau_{i}/L$ Derrida (1983); Akimoto et al. (2016, 2018). When the mean waiting time $\Braket{\tau}$ is finite, we have $\bar{\tau}\rightarrow\Braket{\tau}$ ($L\rightarrow\infty$) by the law of large numbers. Hence, in the large-$L$ limit, $\kappa(1)$ does not depend on the energy landscape. When the mean waiting time diverges (e.g., the power law distribution (38) for $\alpha<1$), the law of large numbers does not hold, and $\kappa(1)$ shows the sample-to-sample fluctuation Akimoto et al. (2016, 2018). In the many-particle system, the current fluctuation $\kappa(N)$ depends on the energy landscape. Thus, $\kappa(N)$ in the intermediate density range exhibits the sample-to-sample fluctuation. As a result of the many-body interactions, the sample-to-sample fluctuation of $\kappa(N)$ emerges in the intermediate density range, even in a one-particle system that does not exhibit the sample-to-sample fluctuation.

### v2

@@ I Introduction
Single file diffusion, where particles diffuse in a narrow channel and cannot pass through each other, is a pedagogical model for interacting many-body systems. It is well known that, on the infinite line, the mean square displacement (MSD) of the tracer particle is given by $\Braket{\delta x(t)^{2}}\propto t^{1/2}$ in the long time limit [1], where $\delta x(t)$ is the displacement of the tracer particle and $\Braket{\cdot}$ is the ensemble average. Various transport phenomena in narrow channels show single file diffusion, for example, molecular diffusion in zeolites [2, 3], DNA-binding proteins [4], transport of confined colloidal particles [5, 6], and transport of water molecules in carbon nanotubes [7].
The symmetric simple exclusion process (SSEP) is a fundamental model of single-file diffusion [8], where particles with hard-core interactions move on a one-dimensional lattice. With its simplicity, this model has led to the derivation of mathematically rigorous results. On the infinite lattice, the large deviation function of the integrated current through the origin has been computed independently by Bethe ansatz [9] and the macroscopic fluctuation theory [10]. Additionally, the large deviation function of the tracer position has been derived [11, 12]. The generalized density profile, which is the correlation between the observation (for example, the tracer position, the integrated current through the origin, and the generalized current) and the density, has also been derived [13, 14, 15]. On the semi-infinite lattice connected to a reservoir, the cumulant generating function of the integrated current has been derived using the macroscopic fluctuation theory [16]. Under the periodic boundary condition, the cumulants of the integrated current have been obtained [17]. In particular, the second moment of the total integrated current $Q(t)$ in time $t$ is derived as a function of the particle density $\rho$:
@@
where $\tau$ is the mean waiting time, and $L$ is the system size [17].
@@
Single-particle diffusion in a heterogeneous environment shows various anomalous behaviors. Such environments are represented by random energy landscapes. The random energy landscape can be categorized into two types: an annealed energy landscape, which changes with time, and a quenched energy landscape, which remains static. The random walk on an annealed energy landscape is often modeled by the continuous-time random walk. In this model, waiting times are independent and identically distributed (IID) random variables, and the MSD exhibits anomalous diffusion when the mean waiting time diverges [18]. The random walk on a quenched energy landscape is called the quenched trap model [19]. In the infinite system, the MSD shows anomalous diffusion when the mean waiting time diverges [19]. In the finite system, observables such as the drift, the diffusion coefficient, and the first passage time exhibit sample-to-sample fluctuations, resulting in non-self-averaging behaviors [20, *AkimotoBarkaiSaito2018, 22, 23, 24].
@@
The influence of heterogeneity on the single-file diffusion has garnered significant attention in the field of statistical mechanics [25, 26, 27, 28, 29]. Recently, the single file diffusion in the system with periodically varying potential and diffusivity has been studied, and the MSD of the tracer particle has been obtained [25]. Specifically, the formula of the MSD is the same as that for the normal single-file diffusion. However, the diffusion coefficient is replaced with the effective diffusion coefficient for a single particle diffusing in the system with periodically varying potential and diffusivity. For the SSEP on an annealed energy landscape, when the mean or variance of waiting time diverges, the MSD is $\Braket{\delta x(t)^{2}}\propto\log^{1/2}t$ or $\Braket{\delta x(t)^{2}}\propto t^{\gamma}$ with $20<\gamma<1/$ 2, respectively [26, 27], i.e., the diffusion in the SSEP on an annealed energy landscape is slower than in the normal SSEP. The impacts of disorder in the asymmetric simple exclusion process (ASEP), where particles in the SSEP are influenced by an external field like an electric field, have been a subject of research for many years [30, 31, 32, 33, 34, 35, 36, *PhysRevE.107.054131, 38, 39, 40, 41, 42]. Under the periodic boundary condition, while, in the ASEP in the homogeneous environment, the relation between current $J$ and particle density $\rho$ is given by $J=\rho(1-\rho)$, in the ASEP on a quenched random energy landscape, it becomes flat around $\rho=1/2$ [30, 31, 32, 33, 34, 35, 36, 37]. Furthermore, the current and the diffusion coefficient exhibit sample-to-sample fluctuations [36, *PhysRevE.107.054131]. While several effects resulting from the interplay between a heterogeneous environment and a many-body system have been uncovered, the study of the SSEP on a quenched random energy landscape remains unexplored. Such heterogeneity is prevalent in various diffusion phenomena, such as the movement of proteins on DNA [43, 44] and water transportation in aquaporin [45] and OmpF [46]. Therefore, understanding current fluctuations in the SSEP on a quenched random energy landscape is crucial.
@@ II Model
We study the SSEP with a defect site to unravel the effects of heterogeneity of the environment. The system consists of $N$ particles on the lattice of $L$ sites with periodic boundary conditions, where each site can accommodate at most one particle. The defect site has a longer mean waiting time compared to other sites, inhibiting particle movement. The mean waiting times at site $i$ are defined as follows:
@@
We explicitly describe the dynamics of particles. Initially, the particles are in an equilibrium ensemble of configurations. The equilibrium ensemble is given by the configuration after particles are randomly placed and diffuse for a long time. In infinite systems, the effects of the initial ensemble are everlasting [47]. At time $t=0$, each particle $i$ is assigned the waiting time $t_{i}$ from $\psi^{(x_{i})}(t_{i})$, where $x_{i}$ is the site number of particle $i$. Once the waiting time $t_{i}$ has elapsed, particle $i$ attempts to jump to the right site with a probability of $1/2$ and to the left site with a probability of $1/2$. The jump is successful only if the target site is unoccupied. Following a successful or unsuccessful jump attempt, particle $i$ is given a new waiting time based on either $\psi^{(x_{i}\pm 1)}(t_{i})$ or $\psi^{(x_{i})}(t_{i})$, respectively.
@@ III.1 Correlation between total integrated current and occupation number
To obtain the correlation between the total integrated current and the occupation number, we derive $\Braket{Q\eta_{2}}_{L-1}$. When $N=L-1$, there is a single hole. The hole at site $i$ jump to the right site with rate $1/(2\tau_{i+1})$ and to the left site with rate $1/(2\tau_{i-1})$. Using the result in Ref. [48], the second moment of the total integrated current for $N=L-1$ is given by
@@ III.2 Second moment of total integrated current
Based on the results of the previous section, the second moment of the total integrated current is represented by
@@
We note that the function $g(N,L,\mu)$ follows Eq. (30), and $g(2,L,\mu)$ and $g(L-2,L,\mu)$ are obtained by the numerical simulations. We compare the theoretical result, Eq. (33), with the numerical result. Figure 1 shows good agreement between numerical simulations and the theory. The density dependence of $\kappa(N)$ for the SSEP with a defect site is different from that for the homogeneous system (1). In particular, it becomes asymmetric around $N/L=1/2$. Furthermore, in the intermediate and high-density regime, $\kappa(N)$ for the SSEP with a defect site is smaller than that for the homogeneous system. This result is that the defect hiders the particle movement, leading to a reduction in fluctuations. On the other hand, in the low-density regime, $\kappa(N)$ for the SSEP with a defect site exceeds that for the homogeneous system. This property is counterintuitive because despite the defect impeding the movement of particles, the fluctuations increase. In terms of diffusivity, adjusting the defect allows for control over the particle’s diffusivity.
@@ IV Application: SSEP on a quenched random energy landscape
Figure 2: Relation between $\kappa(N)$ and density, corresponding to the SSEP on a quenched random energy landscape ($L=100$, $T=1$, $\tau_{c}=1$): (a)-(d) $\Braket{E}<T$ and (e)-(h) $\Braket{E}>T$. (a),(e) The energy depth distribution is the exponential distribution ($T/T_{g}=1.5$ for (a) and $T/T_{g}=0.5$ for (e)). (b),(f) The energy depth distribution is the gamma distribution ($T/T_{g}=2.5,~{}k=2$ for (b) and $T/T_{g}=2.5,~{}k=4$ for (f)). (c),(g) The energy depth distribution is the half-gaussian distribution ($\sigma^{2}=1$ for (c) and $\sigma^{2}=4$ for (g)). (d),(h) The energy depth distribution is the uniform distribution ($\varepsilon=1$ for (d) and $\varepsilon=3$ for (h)). Circles are the numerical simulation of the dynamics of the SSEP on a quenched random energy landscape. Solid lines represent Eq. (43), where we use $g(2,L,\mu)$ and $g(L-2,L,\mu)$ obtained by the numerical simulations of the dynamics of the SSEP with a defect site. Dashed lines show the result in the homogeneous system, Eq. (1), where $\tau$ is set to equal to the average of the waiting time of the SSEP on the quenched random energy landscape, i.e., $\tau=\sum_{i=1}^{L}\tau_{i}/L$.
Here, we apply our theory of the current fluctuations of the SSEP with a defect site to the SSEP on a quenched random energy landscape. Quenched disorder means that when realizing the random energy landscape, it does not change with time. The quenched random energy landscape for the SSEP assigns random energy traps to sites, influencing particle jump rates. More precisely, at each site, the depth $E>0$ of the energy trap is randomly assigned. The energies of the traps are IID random variables following distribution $\phi(E)$. A particle escapes from the trap and attempts to jump to the nearest neighbors. Waiting times at site $i$ are IID random variables with an exponential distribution, $\psi^{(i)}(\tau)=\tau_{i}^{-1}\exp(-\tau/\tau_{i})$. The mean waiting time follows the Arrhenius law, $\tau_{i}=\tau_{c}\exp(E_{i}/T)$, where $E_{i}$ is the depth of the energy at site $i$, $T$ the temperature, and $\tau_{c}$ typical time.
@@
Based on the Arrhenius law, the probability that the mean waiting time $\tau$ is smaller than $x$ is given by $\Pr(\tau\leq x)\cong\Pr(E\leq T\ln(x/\tau_{c}))$. Hence, the distributions of the mean waiting times follow
@@
$\psi(\tau)\cong\frac{\alpha^{k}\tau_{c}^{\alpha}}{\Gamma(k)}\left[\ln\left(\frac{\tau}{\tau_{c}}\right)\right]^{k-1}\tau^{-(1+\alpha)}\quad(\tau\geq\tau_{c})$ (39)
@@
We compare the numerical results with the theoretical result (1) in the homogeneous environment [see Fig. 2]. In the intermediate and high-density regime, $\kappa(N)$ for the SSEP on the quenched random energy landscape is consistently smaller than that for the homogeneous environment. In the low-density regime, $\kappa(N)$ for the SSEP on the quenched random energy landscape is approximately equal to that for the homogeneous environment [Figs. 2(a)-(d), (f)-(h)] or smaller than it [Fig. 2(e)]. $\kappa(N)$ for the SSEP on the quenched random energy landscape does not exceed that in the homogeneous environment because the heterogeneity in the quenched random energy landscape is stronger than that in the system with a defect site.
@@ V Discussion
We compare the current fluctuation in the single-particle system and the many-particle system. In the single-particle system, the current fluctuation is the same as the mean square displacement, i.e., $\kappa(1)=\bar{\tau}^{-1}$, where $\bar{\tau}=\sum_{i=1}^{L}\tau_{i}/L$ [48, 20, 21]. When the mean waiting time $\Braket{\tau}$ is finite, we have $\bar{\tau}\rightarrow\Braket{\tau}$ ($L\rightarrow\infty$) by the law of large numbers. Hence, in the large-$L$ limit, $\kappa(1)$ does not depend on the energy landscape. When the mean waiting time diverges (e.g., the power law distribution (38) for $\alpha<1$), the law of large numbers does not hold, and $\kappa(1)$ shows the sample-to-sample fluctuation [20, 21]. In the many-particle system, the current fluctuation $\kappa(N)$ depends on the energy landscape. Thus, $\kappa(N)$ in the intermediate density range exhibits the sample-to-sample fluctuation. As a result of the many-body interactions, the sample-to-sample fluctuation of $\kappa(N)$ emerges in the intermediate density range, even in a one-particle system that does not exhibit the sample-to-sample fluctuation.
Figure 4: Comparison between quenched and annealed results of the $\kappa(N)$-density relation ($L=100$, $T=1$, $\tau_{c}=1$). The distribution of the energy depth is the exponential distribution ($T/T_{g}=2.5$ for (a) and $T/T_{g}=1.5$ for (b)). Squares are the results of the numerical simulation of the dynamics of the SSEP on a quenched random energy landscape. Circles are the results of the numerical simulation of the dynamics of the SSEP on an annealed random energy landscape.
We compare the current fluctuations in the SSEP on the quenched and annealed random energy landscape. The distinction between these two types of landscapes lies in whether the waiting time distribution depends on the site. For the quenched random energy landscape, the waiting time distribution depends on the site, while for the annealed random energy landscape, it does not. For the ASEP, the current-density relation exhibits differences [41, 36]. In particular, for the quenched random energy landscape, the relation becomes flat around $N/L=1/2$. In contrast, for the annealed random energy landscape, the relation becomes asymmetric around $N/L=1/2$ when the variance of waiting time diverges. To distinguish between the cases where the variance of the waiting time is finite and where it diverges, we consider that the distribution of the energy depth follows the exponential distribution. For the annealed random energy landscape, the waiting time distribution follows the power-law distribution. Figure 4 shows that the relation between $\kappa(N)$ and density becomes asymmetric around $N/L=1/2$. When the variance is finite [Fig. 4(a)], $\kappa(N)$ for the annealed random energy landscap is greater than that for the quenched random energy landscap. Unlike the case of a finite variance, $\kappa(N)$ for the annealed random energy landscape is suppressed more than that for the quenched random energy landscape [Fig. 4(b)]. The asymmetry observed for the annealed disorder arises because, similar to the ASEP [41], the interactions with surrounding particles lead to the divergence of the average waiting time. Consequently, $\kappa(N)$ is surpressed significantly.
We discuss how to control the current fluctuations. In the homogeneous environment, we can control the current fluctuations by adjusting the particle density. On the other hand, in the heterogeneous environment, we can control the current fluctuations by regulating not only the particle density but also the shape of the energy landscape. For example, in the drug delivery device with uniform radii pores, the radii of the pores play a crucial role in controlling the drug release rate [49]. Extending this concept, we propose that controlling current fluctuations in heterogeneous environments could be applied to drug delivery devices with non-uniform pore radii, where the shape of the energy landscape corresponds to the variation in pore radii. Additionally, the understanding of controlling current fluctuations in heterogeneous environments may provide insights into biological transport phenomena, such as DNA-binding proteins [43, 44, 4] and biological channels [45, 46].